lisa spent $44 on shirts.tee shirts cost $10 and long sleeves shirts cost $4. if she bougt a total of 8 then how many of each kind did she buy?
Answer by checkley77(12569)

44=10X+4(8-X)
44=10X+32-4X
6X=44-32
6X=12
X=12/6
X=2 NUMBER OF LONG SLEEVE SHIRTS
8-2=6 SHORT SHIRTS.
PROOF:
44=12*2+5*4
44=24+20
44=44